OpenStudy (anonymous):

sin^-1tan(pi/4) the key is pi/2 Please! explain how get the answer,ty

OpenStudy (anonymous):

\[\sin^-1(\tan\frac{\pi}{4})\] the key is pi/2

OpenStudy (anonymous):

tan(pi/4) = 1 So sin^-1 of 1 = pi/2

OpenStudy (anonymous):

use unit circle

OpenStudy (lgbasallote):

sin^-1[tan pi/4] = sin^-1[tan45] = sin^-1[1] = 90 or pi/2